# Locating Superspreaders in SI Cascades

### 05 Mar 2024 v1

@@ 4 Cascade dynamics
$\operatorname{\mathbb{P}}\left(\left.v\in I[0,t+\epsilon]\right|I[0,t]\right)=\epsilon\lambda|{\mathcal{N}}(v)\cap I[0,t]|+o(\epsilon).$ (4.1)
For a given vertex $v\in V$, we formally define its infection time in the SI model to be
$T(v):=\inf\{t\geq 0:v\in I[0,t]\}.$
It is well-known that the dynamics (4.1) is equivalent to First Passage Percolation on $G$ with $\mathrm{Exp}(\lambda)$ edge weights, which leads to an extremely useful representation of the SI process.
@@
In a bit more detail, let $\{{\mathcal{F}}_{s}\}_{s\geq 0}$ denote the natural filtration corresponding to the SI process. Fix $t\geq 0$, and let $\{F(e)\}_{e\in E}$ be a collection of i.i.d. $\mathrm{Exp}(\lambda)$ random variables that are independent of ${\mathcal{F}}_{t}$. For a path [^1] Recall that a path is a finite sequence of edges $e_{1},\ldots,e_{m}$ such that for each $i\geq 1$, $e_{i}$ and $e_{i+1}$ share exactly one endpoint. $P$ in the graph, we define the weight of the path to be $\mathsf{weight}(P):=\sum_{e\in P}F(e)$. Then we have the following result.
@@
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Let ${\mathcal{P}}(v,t)$ be the collection of paths from $I[0,t]$ to $v$ where only the first edge in the path has an endpoint in $I[0,t]$. Then the conditional distribution of $T(v)$ with respect to ${\mathcal{F}}_{t}$ is given by
@@
This result is standard, and essentially follows from the memoryless property of exponential distributions. For completeness, we provide the proof in Appendix A. [say some further remarks about FPP, this is a bit abrupt]
Using the representation in Proposition 4.1, we can derive a number of useful results on the infection times of vertices.
@@
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Then $T(v)-t$ is stochastically dominated by a $\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)$ random variable.
@@
$T(v)-t$ $=\min_{P\in{\mathcal{P}}(v,t)}\mathsf{weight}(P)$
$\leq\min_{P\in{\mathcal{P}}(v,t):|P|=1}\mathsf{weight}(P)$
$=\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v).$
@@
$\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\sim\mathrm{Exp}\left(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right),$
@@
Suppose that $G$ is a tree. Let $t\geq 0$, condition on ${\mathcal{F}}_{t}$, and suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Then $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda)$ random variable.
Proof.
Since $G$ is a tree, there is a unique path from the infection source $s$ to $v$. As a result, $v$ can never have more than one infected neighbor, so $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda)$ random variable. ∎
[lower bound on infection time]
Lemma 4.4.
Let $t\geq 0$, condition on ${\mathcal{F}}_{t}$, and suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Then $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda|{\mathcal{N}}(v)|)$ random variable.
@@
[number of infected neighbors of high-deg vertex by the time it gets infected]
We next define a useful event concerning the number of infected neighbors of high-degree vertices.
[potentially change to $n^{1/2+\epsilon}$, or similar. It would look cosmetically simpler and probs would be easier to digest.]
@@
$|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,T(v)]|\leq 2d\sqrt{\deg(v)}\log^{2}n,$
for all $v\in V$ such that $\deg(v)\geq D$.
@@
It holds that $\operatorname{\mathbb{P}}({\mathcal{A}})=1-o(1)$.
@@
Let $\deg(v)\geq D$, and define the event
${\mathcal{E}}_{v}:=\left\{|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,T(v)]|\geq 2d\sqrt{\deg(v)}\log^{2}n\right\}.$
Let $T$ be the (random) time at which $\sqrt{\deg(v)}\log n$ neighbors of $v$ are first infected.
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Let us also assume that $T=t$ (an event that is ${\mathcal{F}}_{t}$-measurable), so that $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|=\sqrt{\deg(v)}\log n$. If $T(v)\leq t$, then we have that $\operatorname{\mathbb{P}}({\mathcal{A}}|{\mathcal{F}}_{t})=0$. Otherwise, Lemma 4.2 implies that $T(v)-t$ is stochastically dominated by a $\mathrm{Exp}(\lambda\sqrt{\deg(v)}\log n)$ random variable. Hence
$\operatorname{\mathbb{P}}\left(\left.T(v)\geq t+\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)\leq\frac{1}{n}.$ (4.2)
Next, we upper bound the number of neighbors of $v$ that become infected in the interval $[t,t+1/(\lambda\sqrt{\deg(v)})]$. By our assumption on the distance between high-degree vertices [ref somewhere], all neighbors of $v$ have degree at most $d$. Lemma 4.4 therefore implies that for $u\in{\mathcal{N}}(v)\setminus{\mathcal{I}}[0,t]$, $T(u)-t$ stochastically dominates a $\mathrm{Exp}(\lambda d)$ random variable, and
$\operatorname{\mathbb{P}}\left(\left.T(u)-t\leq\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)\leq 1-e^{-d/\sqrt{\deg(v)}}\leq\frac{d}{\sqrt{\deg(v)}}.$
It follows that
$\operatorname{\mathbb{E}}\left[\left.|{\mathcal{N}}(v)\cap{\mathcal{I}}[t,T(v)]|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right)\right|{\mathcal{F}}_{t}\right]$ $\leq\operatorname{\mathbb{E}}\left[\left.\left|{\mathcal{N}}(v)\cap{\mathcal{I}}\left[t,t+\frac{1}{\lambda\sqrt{\deg(v)}}\right]\right|\right|{\mathcal{F}}_{t}\right]$
$=\sum_{u\in{\mathcal{N}}(v)\setminus{\mathcal{I}}[0,t]}\operatorname{\mathbb{P}}\left(\left.T(u)\leq t+\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)$
$\leq d\sqrt{\deg(v)}.$
As a consequence,
$\operatorname{\mathbb{E}}\left[\left.\left|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,T(v)]\right|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right)\right|{\mathcal{F}}_{t}\right]$
$\hskip 28.45274pt\leq|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|+\operatorname{\mathbb{E}}\left[\left.|{\mathcal{N}}(v)\cap{\mathcal{I}}[t,T(v)]|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right)\right|{\mathcal{F}}_{t}\right]$
$\hskip 28.45274pt\leq\sqrt{\deg(v)}\log n+d\sqrt{\deg(v)}\leq 2d\sqrt{\deg(v)}\log n,$
@@
$\operatorname{\mathbb{P}}\left(\left.{\mathcal{E}}_{v}\cap\left\{T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right\}\right|{\mathcal{F}}_{t}\right)\leq\frac{1}{\log^{2}n}.$
A union bound along with (4.2) shows that
$\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{t})\leq\operatorname{\mathbb{P}}\left(\left.{\mathcal{E}}_{v}\cap\left\{T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right\}\right|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}\left(\left.T(v)\geq t+\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)\leq\frac{2}{\log^{2}n}.$ (4.3)
@@
which, in light of (4.3), implies that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{T})\leq 2/\log^{2}(n)$. It follows that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=\operatorname{\mathbb{E}}[\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{T})]\leq 2/\log^{2}(n)$. Finally, we note that ${\mathcal{A}}^{c}:=\bigcup_{v\in G:\deg(v)\geq D}{\mathcal{E}}_{v}$, hence a union bound shows that
$\operatorname{\mathbb{P}}({\mathcal{A}}^{c})\leq\sum_{v\in G:\deg(v)\geq D}\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=o(1).$

### 06 Mar 2024 v2

# Finding Super-spreaders in Network Cascades

@@ 2.1 Estimating high-degree vertices when $\alpha>3/4$
$J(t):=\operatorname{\mathbb{E}}\left[\left.\frac{d}{dt}I(t)\right|{\mathcal{F}}_{t}\right]=\lambda\sum_{v\in V\setminus{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|=\lambda\operatorname{\mathsf{cut}}({\mathcal{I}}(t)).$
Now suppose that a vertex $v$ joins the cascade at time $t$ (that is, $T(v)=t$). Since no other vertices can be infected at the exact same time due to the continuous-time nature of the cascade, we have
$J(t)-J(t^{-})=\lambda\left(|{\mathcal{N}}(v)\setminus{\mathcal{I}}(t)|-|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\right)=\lambda\deg(v)-2\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|.$ (2.1)
The first equality in (2.1) can be explained as follows. When $v$ joins the cascade, all of its susceptible neighbors, i.e., those in ${\mathcal{N}}(v)\setminus{\mathcal{I}}(t)$, become exposed. On the other hand, the edges which connect $v$ to ${\mathcal{I}}(t)$ – namely, edges which connect $v$ to ${\mathcal{N}}(v)\cap{\mathcal{I}}(t)$ – can no longer lead to infection events, hence the contribution for these edges is subtracted. See Figure 2 for a visualization of this idea.
Crucially, (2.1) shows that as long as a vertex is infected before most of its neighborhood (i.e., $|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|$ is much smaller than $\deg(v)$), the second-order changes in the infection curve, captured by $J(t)-J(t^{-})$, can reveal the degree of a vertex. However, since $J(t)$ is not a directly observable quantity, we must approximate it from the cascade traces. A natural way to do so is by choosing $\delta>0$ and using the discrete derivative $I[t,t+\delta]/\delta$ as a proxy for $J(t)$. Similarly, we can use $I[t-\delta,t]/\delta$ as a proxy for $J(t^{-})$. We can therefore approximate the degree of $v$ using
$\widehat{\deg}_{\delta}(v):=\frac{I[T(v),T(v)+\delta]-I[T(v)-\delta,T(v)]}{\delta}.$
In words, we estimate the degree of $v$ by computing the discrete second derivative of the infection curve when $v$ gets infected. As shown through the empirical example in Figure 3, this statistic is quite effective in identifying the infection time of a high-degree vertex from just a single cascade.
Figure 3: Plots of the discrete first $(a)$ and second derivative $(b)$ of the infection curve with $\delta=0.075$ generated from a graph $G$ with approximately 500,000 vertices and one high-degree vertex. The infection time of the high-degree vertex can be identified by the red dotted line in $(a)$ and the large peak in the second derivative plot $(b)$, highlighted by the red rectangle. $G$ was chosen to be a balanced, 5-regular tree of height 8, where one vertex in the 6th layer of the tree has degree 7500.
In light of (2.1), our estimator for $\mathsf{highdeg}(G)$ returns the set of all vertices $v$ for which $\widehat{\deg}_{\delta}(v)$ is large in most cascade traces; see Algorithm 3. As the Theorem 2.1 shows, this method succeeds with high probability provided $\alpha>3/4$.
  Algorithm 1 Finding high-degree vertices via second derivative thresholding
Δ=Δ1:Cascade traces $\mathbf{T}_{1},\ldots,\mathbf{T}_{K}$ and parameters $\delta,\tau>0$
Δ=Δ2:A set $\widehat{\mathrm{HD}}\subset V$
Δ=Δ3:For each $i\in[K]$ and $v\in V$ compute
$\widehat{\deg}_{\delta,i}(v):=\frac{I_{i}[T_{i}(v),T_{i}(v)+\delta]-I_{i}[T_{i}(v)-\delta,T_{i}(v)]}{\delta}.$
Δ=Δ4:$\widehat{\mathrm{HD}}\leftarrow\emptyset$.
Δ=Δ5:For each $v\in V$: If $\widehat{\deg}_{\delta,i}(v)\geq\tau$ for at least $K/2$ cascades, then $\widehat{\mathrm{HD}}\leftarrow\widehat{\mathrm{HD}}\cup\{v\}$.
Δ=Δ6:Return $\widehat{\mathrm{HD}}$.
Theorem 2.1.
Suppose that $\alpha>3/4$ and that
$\delta=\frac{1}{n^{\alpha-1/4}},\hskip 28.45274pt\tau=\frac{n^{\alpha}}{\log n},\hskip 28.45274ptK\geq\frac{15}{\alpha-3/4}.$
Additionally assume that $d=n^{o(1)}$ and $D=n^{\alpha}$. Then for any $G\in{\mathcal{G}}(n,m,d,D)$, the output of Algorithm 3 is equal to $\mathsf{highdeg}(G)$, with probability $1-o(1)$.
Remark 2.2.
A useful feature of Algorithm 3 is that it is adaptive to the unknown parameters $m$ (the number of high-degree vertices) and $\lambda$ (the spreading rate of the cascade). In contrast, the prior work on structure learning leverages likelihood-based methods which require precise knowledge of the cascade model [39, 1].
We note that Theorem 1.3 follows immediately from Theorem 2.1 as a corollary. We provide a sketch of the proof of Theorem 2.1 below, and defer the full details to Sections 4 and 5.
Proof sketch of Theorem 2.1.
Let $\gamma<\alpha$ be sufficiently close to $\alpha$. The strategy of our proof is to show, for all $v\in V$, that
$\left|\widehat{\deg}_{\delta,i}(v)-\lambda\deg(v)\right|\leq n^{\gamma},$ (2.2)
for at least half of the cascades. If (2.2) is true for a particular $i\in[K]$ and $v$ is a low-degree vertex, then $\widehat{\deg}_{\delta,i}(v)\leq\lambda\deg(v)+n^{\gamma}\leq 2n^{\gamma}$. Else if $v$ is high-degree, then $\widehat{\deg}_{\delta,i}(v)\geq\lambda\deg(v)-n^{\gamma}\geq\frac{\lambda}{2}n^{\alpha}$. If $\lambda$ is not known a priori, any choice of threshold satisfying $\tau=n^{\alpha-o(1)}$ works in distinguishing the two cases. In particular, $\tau=n^{\alpha}/\log(n)$ suffices.
Due to the way $\widehat{\deg}_{\delta,i}(v)$ is computed, establishing (2.2) reduces to showing that $I[t,t+\delta]/\delta$ concentrates around $J(t)$ and $I[t,t-\delta]/\delta$ concentrates around $J(t^{-})$ for the values of $t$ corresponding to vertices’ infection times. For this task, choosing the right value of $\delta$ is crucial. If $\delta$ is too small, then $I[t,t+\delta]$ may also be quite small, making meaningful concentration impossible. On the other hand, if $\delta$ is too large, then there may be a large number of low-degree vertices that consistently infected around the same time as a high-degree vertex. If $u$ is one such low-degree vertex, then $\widehat{\deg}_{\delta}(u)$ can be quite large across multiple traces, leading to false positives.
Examining these conditions in more detail exposes why we must assume $\alpha>3/4$ for Algorithm 3 to succeed. Let us first consider the behavior of $I[t,t+\delta]/\delta$ (the analysis of $I[t-\delta,t]/\delta$ is similar). Since $I[t,t+\delta]$ is derived from a counting process, we can deduce a subgaussian-type concentration of $I[t,t+\delta]$ around $\delta J(t)$ (see Lemma 5.3 in Section 5.1). It follows that
$I[t,t+\delta]=\delta J(t)\pm O\left(\sqrt{\delta J(t)}\right),$ (2.3)
where the second term represents the random fluctuations around $J(t)$. A direct consequence of (2.3) is that the deviation of $I[t,t+\delta/\delta]/\delta$ from $J(t)$ is at most $O(\sqrt{J(t)/\delta})$. In order to correctly estimate high-degree vertices through (2.1), this deviation needs to be smaller than $D=n^{\alpha}$. Since all but a $o(1)$ fraction of vertices in $G$ has degree at most $d=n^{o(1)}$, it follows from (2.1) that $J(t)\leq n^{1+o(1)}$. Hence we require that $\sqrt{n^{1+o(1)}/\delta}=o(n^{\alpha})$, which is equivalent to
$\delta=\omega\left(n^{-(2\alpha-1)-o(1)}\right).$ (2.4)
Next, we need to ensure that there is no low-degree vertex $u$ that is infected around the same time as a high-degree vertex in at least half the cascade traces. Otherwise, as we mentioned earlier, $\widehat{\deg}_{\delta,i}(u)$ could be quite large, which would imply that the output of Algorithm 3 contains false positives. With this scenario in mind, suppose that a low-degree $u$ becomes infected first at time $t$, and that point in time, a high-degree $v$ has $M$ infected neighbors. By (1.1), it holds for small $\delta$ that
$\operatorname{\mathbb{P}}(|T(u)-T(v)|\leq\delta|{\mathcal{F}}_{t})=\operatorname{\mathbb{P}}(v\in{\mathcal{I}}(t+\delta)|{\mathcal{F}}_{t})=\delta M+o(\delta).$
The probability that $u$ and $v$ are close in at least half of the cascade traces is therefore at most $O((\delta M)^{K/2})$. A union bound over pairs of vertices shows that
$\operatorname{\mathbb{P}}(\exists u,v\in V:|T_{i}(u)-T_{i}(v)|\leq\delta\text{in at least half the cascades})\leq O\left(n^{2}(\delta M)^{K/2}\right).$ (2.5)
To show that the right hand side of (2.5) is $o(1)$, we first need to choose $\delta$ such that $\delta M=o(1)$. Although $M$ can in principle be as large as $\deg(v)$ (i.e., if all of $v$’s neighbors are infected before $v$), we show that with high probability $M\leq\sqrt{n}$ (see Lemma 3.7 in Section 3). Hence we may choose $\delta$ satisfying
$\delta=o\left(n^{-1/2}\right).$ (2.6)
Writing $\delta=n^{-\beta}$ for $\beta>0$, (2.4) and (2.6) imply that $\beta\in(1/2,2\alpha-1)$. This interval is nonempty when $\alpha>3/4$, and we may choose $\beta:=\alpha-1/4$ to satisfy (2.4) and (2.6). Finally, since $\delta=n^{-(\alpha-1/4)}$ and $M\leq\sqrt{n}$, $K\geq\Omega((\alpha-3/4)^{-1})$ is needed for the right hand side of (2.5) to be $o(1)$. ∎
2.2 Impossibility of estimation when $\alpha\in(0,1/2)$
Connections to sparse mixture detection.
The cornerstone of our proof is a correspondence between high-degree estimation and the problem of sparse mixture detection. In a bit more detail, let $G^{\prime}$ be a graph on $n/2$ vertices of maximum degree $n^{o(1)}$ (i.e., all its vertices are low-degree). The graph $G$ will be formed by adding an additional $n/2$ vertices labelled $21,\ldots,n/$ 2 to $G^{\prime}$, along with a single edge per added vertex which connects to a vertex in $G^{\prime}$. A natural way of generating a graph $G$ without any high-degree vertices is to let each of the new $n/2$ vertices connect to a uniform random parent in $G^{\prime}$. Through standard probabilistic arguments, it can be seen that no more than $O(\log n)$ neighbors are added to any particular vertex in $G^{\prime}$, so $G$ contains no high-degree vertices.
A simple modification of this procedure can also generate graphs with a high-degree vertex. Suppose we fix a vertex $v$ in $G^{\prime}$. For each of the additional $n/2$ vertices, with probability $2D/n$, it connects to $v$. Otherwise, with probability $1-2D/n$, it connects to a uniform random vertex in $G^{\prime}$. In expectation, $v$ receives $D$ additional connections (so, in particular, $\mathrm{deg}(v)\geq D$ in expectation). As before, all of the other vertices receive at most $O(\log n)$ additional neighbors, with high probability. Hence we obtain a graph $G$ in which $v$ is the only high-degree vertex.
Observe that in the second procedure, the distribution of the infection times of each of the added $n/2$ vertices can be viewed as a mixture of two distributions. To be precise, let $P$ be the distribution of a vertex’s infection times in the observed cascade traces if they connect to a uniform random neighbor in $G^{\prime}$. Similarly, let $Q$ be the distribution of a vertex’s infection times if they connect to $v$. Then the task of differentiating between the two generative models for $G$ – or equivalently, of testing whether $v$ is high-degree – can be phrased as the following hypothesis testing problem:
$H_{0}$ $:\mathbf{T}(1),\ldots\mathbf{T}(n/2)\sim P\hskip 28.45274ptH_{1}:\mathbf{T}(1),\ldots,\mathbf{T}(n/2)\sim\left(\left(1-\epsilon_{n}\right)P+\epsilon_{n}Q\right),$ (2.7)
where we set $\epsilon_{n}:=2D/n$ above, and we denote $\mathbf{T}(j):=(T_{1}(j),\ldots,T_{K}(j))$ where $T_{i}(j)$ is the infection time of vertex $j$ in the $i$th cascade.
It turns out that (2.7) is a special case of a well-studied problem known as sparse mixture detection, which dates back to the work of Dobrusin [17]. Burnashev [12] studied a variant related to Gaussian processes, and Ingster [28] identified a subtle phase transition between the possibility and impossibility of detection for mixtures of Gaussian distributions. The problem was popularized by the work of Donoho and Jin [18], who formulated an adaptive method in place of the likelihood ratio called higher criticism, again for mixtures of Gaussians. Later, Cai and Wu [14] developed an analysis of the sparse mixture problem for generic distributions. We remark that Mossel and Roch [38] used sparse mixture detection in the inference of combinatorial structure, though in a very different context.
Impossibility of detecting high-degree vertices.
We solve our instance of the sparse mixture detection problem (2.7) using the techniques developed by Cai and Wu [14]. That is, we bound the total variation distance between the two distributions in (2.7) by the $\chi^{2}$ divergence since it tensorizes and because it is easier to study for problems involving mixtures. Our analysis shows that if $K$ satisfies (1.3), then the $\chi^{2}$ divergence tends to zero. As a result, it is impossible to distinguish between scenarios where there are no high-degree vertices ($H_{0}$) and where $v$ is a high-degree vertex ($H_{1}$).
Putting everything together.
Finally, we show that the impossibility of detection implies that no estimator can successfully output the correct high-degree vertex with probability greater than $o(1)$. We remark that the ensemble ${\mathcal{H}}$ and distribution $\mu$ stated in the theorem essentially corresponds to the random constructions for $G$ described in this subsection. For more details, see Section 6.
3 Preliminaries: Cascade dynamics and infection times
In this section we derive several useful results on the behavior of the SI process (1.1) and the distribution of infection times. Throughout, we will assume that ${\mathcal{I}}(t)$ is a realization of $\mathcal{SI}(G,v_{0})$ for a particular graph $G\in{\mathcal{G}}(n,m,d,D)$ and vertex $v_{0}\in V$.
@@ 3 Preliminaries: Cascade dynamics and infection times
We begin by discussing an important equivalence between (1.1) and First Passage Percolation on $G$ with $\mathrm{Exp}(\lambda)$ edge weights. In a bit more detail, let $\{{\mathcal{F}}_{t}\}_{t\geq 0}$ denote the natural filtration corresponding to the SI process. Fix $t\geq 0$, and let $\{F(e)\}_{e\in E}$ be a collection of i.i.d. $\mathrm{Exp}(\lambda)$ random variables that are independent of ${\mathcal{F}}_{t}$. For a path [^1] Recall that a path is a finite sequence of edges $e_{1},\ldots,e_{m}$ such that for each $i\geq 1$, $e_{i}$ and $e_{i+1}$ share exactly one endpoint. $P$ in the graph, we define the weight of the path to be $\mathsf{weight}(P):=\sum_{e\in P}F(e)$. Then we have the following result.
@@
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Suppose that $v\in V\setminus{\mathcal{I}}(t)$. Let ${\mathcal{P}}(v,t)$ be the collection of paths from ${\mathcal{I}}(t)$ to $v$ where only the first edge in the path has an endpoint in ${\mathcal{I}}(t)$. Then the conditional distribution of $T(v)$ with respect to ${\mathcal{F}}_{t}$ is given by
@@
This result is standard, and essentially follows from the memoryless property of exponential distributions. We defer the reader to [5, Chapter 6], or to Section 7 for full details of the proof.
Using the representation in Proposition 3.1, we can derive a number of useful results on the infection times of vertices. We start by establishing probabilistic upper and lower bounds on conditional infection times of vertices.
@@
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Suppose that $v\in V\setminus{\mathcal{I}}(t)$. Then $T(v)-t$ is stochastically dominated by a $\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|)$ random variable.
@@
$T(v)-t$ $=\min_{P\in{\mathcal{P}}(v,t)}\mathsf{weight}(P)\leq\min_{P\in{\mathcal{P}}(v,t):|P|=1}\mathsf{weight}(P)=\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)}F(u,v).$
@@
$\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}(t)}F(u,v)\sim\mathrm{Exp}\left(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|\right),$
@@
Let $t\geq 0$, condition on ${\mathcal{F}}_{t}$, and suppose that $v\in V\setminus{\mathcal{I}}(t)$. Then $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda|{\mathcal{N}}(v)|)$ random variable.
@@
We next define a useful event concerning the number of infected neighbors of high-degree vertices. We remark that the control of this quantity is especially important for us, as it is equal to the bias of our estimate of vertex degrees (see (2.1)).
@@
$|{\mathcal{N}}(v)\cap{\mathcal{I}}(T(v))|\leq 2d\sqrt{n}\log^{2}n,$
for all $v\in V$.
Remark 3.5.
As we assume throughout this paper that $d=n^{o(1)}$, we will often use that $|{\mathcal{N}}(v)\cap{\mathcal{I}}(T(v))|\leq n^{1/2+o(1)}$ when using the event ${\mathcal{A}}$.
Remark 3.6.
Assuming that $G\in{\mathcal{G}}(n,m,d,D)$, the event ${\mathcal{A}}$ is only relevant for high-degree vertices, since
$|{\mathcal{N}}(v)\cap{\mathcal{I}}(T(v))|\leq\deg(v)=n^{o(1)}\leq 2d\sqrt{n}\log^{2}n,$
otherwise.
@@
Suppose that $m=o(\log^{2}n)$. Then $\operatorname{\mathbb{P}}({\mathcal{A}})=1-o(1)$.
@@
If $\deg(v)\leq 2d\sqrt{n}\log^{2}n$, then
$|{\mathcal{N}}(v)\cap{\mathcal{I}}(T(v))|\leq|{\mathcal{N}}(v)|\leq 2d\sqrt{n}\log^{2}n,$
which satisfies the condition in ${\mathcal{A}}$ corresponding to $v$. It therefore suffices to consider the case $\deg(v)\geq 2d\sqrt{n}\log^{2}n$. Define the event
${\mathcal{E}}_{v}:=\left\{|{\mathcal{N}}(v)\cap{\mathcal{I}}(T(v))|\geq 2d\sqrt{n}\log^{2}n\right\},$
and let $T$ be the stopping time at which $\sqrt{n}\log n$ neighbors of $v$ are first infected.
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Let us also assume that $T=t$ (an event that is ${\mathcal{F}}_{t}$-measurable), so that $|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|=\sqrt{n}\log n$. If $T(v)\leq t$, then we have that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{t})=0$. Otherwise, Lemma 3.2 implies that $T(v)-t$ is stochastically dominated by a $\mathrm{Exp}(\lambda\sqrt{n}\log n)$ random variable. Hence
$\operatorname{\mathbb{P}}\left(\left.T(v)\geq t+\frac{1}{\lambda\sqrt{n}}\right|{\mathcal{F}}_{t}\right)\leq\frac{1}{n}.$ (3.1)
Next, we upper bound the number of neighbors of $v$ that become infected in the interval $[t,t+1/(\lambda\sqrt{n})]$. Since no two high-degree vertices are neighbors for $G\in{\mathcal{G}}(n,m,d,D)$ (see Definition 1.1), all neighbors of $v$ have degree at most $d$. It follows that for $u\in{\mathcal{N}}(v)\setminus{\mathcal{I}}(t)$, we have that
($\operatorname{\mathbb{P}}\left(\left.\{T(u)\leq T(v)\}\cap\left\{T(v)<t+\frac{1}{\lambda\sqrt{n}}\right\}\right|{\mathcal{F}}_{t}\right)$ $\leq\operatorname{\mathbb{P}}\left(\left.T(u)<t+\frac{1}{\lambda\sqrt{n}}\right|{\mathcal{F}}_{t}\right)$
$\leq 1-e^{-d/\sqrt{n}}\leq\frac{d}{\sqrt{n}}.$
Above, the first inequality on the second line follows since $T(u)-t$ stochastically dominates a $\mathrm{Exp}(\lambda d)$ random variable, in light of Lemma 3.3. Next, summing over all $u\in{\mathcal{N}}(v)\setminus{\mathcal{I}}(t)$ shows that
$\operatorname{\mathbb{E}}\left[\left.|({\mathcal{N}}(v)\cap{\mathcal{I}}(T(v)))\setminus{\mathcal{I}}(t)|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{n}}\right)\right|{\mathcal{F}}_{t}\right]\leq\frac{d|{\mathcal{N}}(v)\setminus{\mathcal{I}}(t)|}{\sqrt{n}}\leq d\sqrt{n},$
where the final inequality uses $|{\mathcal{N}}(v)|\leq n$. As a consequence of the previous display, we have that
$\operatorname{\mathbb{E}}\left[\left.\left|{\mathcal{N}}(v)\cap{\mathcal{I}}(T(v))\right|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{n}}\right)\right|{\mathcal{F}}_{t}\right]$
$\hskip 28.45274pt\leq|{\mathcal{N}}(v)\cap{\mathcal{I}}(t)|+\operatorname{\mathbb{E}}\left[\left.|({\mathcal{N}}(v)\cap{\mathcal{I}}(T(v)))\setminus{\mathcal{I}}(t)|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{n}}\right)\right|{\mathcal{F}}_{t}\right]$
$\hskip 28.45274pt\leq\sqrt{n}\log n+d\sqrt{n}\leq 2d\sqrt{n}\log n,$
@@
$\operatorname{\mathbb{P}}\left(\left.{\mathcal{E}}_{v}\cap\left\{T(v)<t+\frac{1}{\lambda\sqrt{n}}\right\}\right|{\mathcal{F}}_{t}\right)\leq\frac{1}{\log n}.$
The display above combined with (3.1) shows that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{t})\leq 1/\log(n)+1/n\leq 2/\log(n)$.
@@
which, in light of the bound $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{t})\leq 2/\log(n)$, implies that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{T})\leq 2/\log(n)$. It follows that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=\operatorname{\mathbb{E}}[\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{T})]\leq 2/\log(n)$. Finally, we note that
${\mathcal{A}}^{c}=\bigcup_{v\in V:\deg(v)\geq 2d\sqrt{n}\log^{2}n}{\mathcal{E}}_{v},$
hence a union bound over these at most $m$ events shows that
$\operatorname{\mathbb{P}}({\mathcal{A}}^{c})\leq\frac{2m}{\log n}=o(1).$
As the following result shows, the event ${\mathcal{A}}$ is crucially used to obtain sharper characterizations of infection times than what Lemma 3.3 can provide, especially for high-degree vertices.